# Towards Black-Box Membership Inference Attack for Diffusion Models

### v1

@@ 6 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment’s setup in previous papers [5, 18]. The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
@@ 6.2 MIA with DDIM Models
We train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2] and Tiny-Imagenet [3] dataset, using the image generation step $T=1000$ and sampling interval $k=100$ (explained in Section 6.4). For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members. The remaining 50% are utilized as nonmembers. We take [28], [5], [18] as our baseline method. We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$. We will discuss the impact of the diffusion step and the average number in Section 6.4.
Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method CIFAR10 CIFAR100 STL10 TINY-IN
Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP AUC ASR TP
Loss [28] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6 0.94 0.87 29.4
SecMI [5] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9 0.97 0.94 48.4
PIA [18] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8 0.98 0.94 73.0
PIAN [18] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5 0.95 0.93 78.4
ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9 0.99 0.96 78.8
@@
For the difference function $D(x,\hat{x})$, following the setup in [5], we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network [9] and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negative value of the probability predicting as member by the neural network: $D(x,\hat{x})=-f_{R}(v)$.
The result is shown in Table 1. Our method achieve high performance, surpassing several baseline algorithms in most setups, and does not require access to the internal structure of the model. This demonstrates that our algorithm is highly effective and robust.
@@ 6.3 MIA with the Stable Diffusion Model
We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications. First, we follow the experiment setup of [5, 18], use the LAION-5B dataset [36] as member and COCO2017-val [23] as non-member. We randomly select 2500 images in each dataset. We test two scenarios: Knowing the ground truth text, which we denote as Laion5 vs COCO; Generating text through BLIP [21], which we denote as Laion5 BLIP vs COCO.
Next, given the inherent content differences between the COCO dataset and the LAION dataset, as discussed in Section 4, we also include a new evaluation dataset. We randomly generated 500 images using DALL-E 3, using prompts derived from the captions associated with each image in LAION-5. We refer to this new dataset as “LAION by DALL-E”. This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set. This evaluation focuses on style shift rather than content shift between the datasets, enabling a more comprehensive evaluation of Stable Diffusion.
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [44] to measure the differences between two images. Additionally, because the image sizes in these datasets are relatively large, the differences between the original and reconstructed images are generally significant. Therefore, besides the basic algorithm, we also propose a simple variant of our algorithm (named ReDiffuse+) based on methods discussed in Section 4, involving altering the image by API twice independently and directly measuring the differences between the two variational images, rather than comparing the variational images to the original images.
Table 2: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1. Again, previous methods require access to U-Net, whereas our methods do not.
Method Laion5 vs COCO Laion5 vs COCO with BLIP Laion5 vs Laion5 by DALL-E
AUC ASR TP AUC ASR TP AUC ASR TP
Loss [28] 0.62 0.61 13.2 0.62 0.62 13.3 0.52 0.55 7.6
SecMI [5] 0.70 0.65 19.2 0.71 0.66 19.8 0.66 0.64 15.2
PIA [18] 0.70 0.66 19.7 0.73 0.68 20.2 0.65 0.62 14.9
PIAN [18] 0.56 0.53 4.8 0.55 0.51 4.4 0.53 0.52 3.4
ReDiffuse 0.79 0.72 14.4 0.80 0.74 14.8 0.74 0.70 15.3
ReDiffuse+ 0.81 0.75 20.6 0.82 0.75 22.1 0.73 0.71 13.6
The results, presented in Table 2, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net. In addition, previous methods perform poorly on the new evaluation dataset, failing to achieve effective membership inference attack detection. In contrast, our methods still achieve robust performance, indicating their ability to distinguish images in the training set even when differences in image content are excluded.
We also plot ROC curves for the Stable Diffusion model trained on Laion5 setups and the DDIM trained on CIFAR-10 in Figure 4. These curves further demonstrate the effectiveness of our algorithm.
Figure 4: The ROC curves of various methods trained on DDIM and Stable Diffusion. Left: Trained DDIM model on CIFAR-10. Right: Trained Stable Diffusion model on the LAION-5B dataset with the ground truth text from LAION-5B caption. Non-members are sampled from COCO. The curves show that our algorithm outperforms the baseline algorithms.
@@ 6.4 Ablation Studies
In this section, we alter some experimental parameters to test the robustness of our algorithm.
@@
We use different average numbers $n$ and test the influence on the results. The results are shown in Figure 5. We observe that in the DDIM setup, due to the smaller image size in the dataset (32x32), there is a certain variance in the images obtained from each reconstruction. Averaging the images from multiple independent samples as the output $\hat{x}$ can further improve accuracy. In the stable diffusion setup, since the image size in the dataset is larger (512x512), the reconstructed images are more stable and not influenced by perturbations at specific coordinates. Therefore, averaging multiple images is not necessary. The figures of ASR are also plotted in Appendix B.
Figure 5: The impact of average numbers. Left: DDIM model on CIFAR-10. Averaging can further improve the performance of our algorithm. Right: Stable Diffusion model on LAION-5. In the latter case, averaging does not improve performance.
@@
We adjust the diffusion step $t$ to examine its impact on the results. The experiments are conducted using the DDIM model trained on CIFAR-10 with diffusion steps for inference, ranging from 10 to 500, and the outcomes are presented in Figure 6. Our findings indicate that as long as a moderate step (50-300) is chosen, the attack performance remains excellent, demonstrating that our algorithm is not sensitive to the choice of $t$. This further underscores the robustness of our algorithm. We also plot the change of diffusion step for other datasets in the Appendix B.
Figure 6: The impact of diffusion steps. We train a DDIM model on the CIFAR-10 dataset and use different diffusion step for inference. We find that high accuracy can be achieved as long as a moderate step number is chosen. This opens up possibilities for practical applications in real-world scenarios.
@@
In DDIM and the internal structure of Stable Diffusion, the model use a set of steps denoted by $\tau_{1},\tau_{2},\ldots,\tau_{T}$. It samples at each of these steps to create the image. The spacing between these steps is referred to as the sampling interval.
We change the sampling interval $k$ and check the influence on the results. As shown in Figure 7, we adjust this parameter for attacks on both DDIM and stable diffusion. We find that our method achieve high AUC values across different sampling intervals, demonstrating that our detection capabilities are not significantly limited by this parameter. Additionally, in Appendix B, we plot the effect of different sampling interval on ASR and find that the impact is minimal.
@@ 7 Conclusion, Limitations and Future Directions
In this work, we introduce a novel black-box membership inference attack method specifically designed for diffusion models. Our approach only requires access to the variation API of the model, bypassing the need for internal network components such as the U-net. This represents a significant advancement for commercial diffusion models, which typically restrict internal access. We demonstrate the effectiveness of our approach across various datasets, showing that it achieves high accuracy in identifying whether an image was included in the training dataset. Our method offers a practical solution for protecting the copyright of artworks.
However, our method has some limitations, particularly the need for a moderate diffusion step $t$ of the variation API. Our algorithm does not achieve high accuracy if the diffusion step is very high. Additionally, it would be desirable if we could expand the moderate size of our newly proposed evaluation dataset.
Future work could aim to derive more robust algorithms that can be used for a wider range of diffusion step. Improving the interpretability of our method and extending it to other generative models are also valuable directions for further research.
@@ Appendix A Datasets, Models and Hyperparameters
We run experiments on DDIM and Stable Diffusion models, using an NVIDIA RTX 6000 graphics card for all experiments.
For DDIM, we follow the training hyperparameters of [5] to train a DDIM model on the CIFAR-10/100 [20], STL10-Unlabeled [2], and Tiny-Imagenet [3] datasets, using an image generation step of $T=1000$ and a sampling interval of $k=100$. The training iterations are set to 800,000. For all the datasets, we randomly select 50
We take [28], [5], and [18] as our baseline methods. We use their codes from the official repositories and apply the optimal hyperparameters from their papers. For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-5 provided by Huggingface, without further fine-tuning or modifications. We follow the experimental setup of [5, 18], using an image generation step of $T=1000$ and a sampling interval of $k=10$. We use the LAION-5B dataset [36] as the member set and COCO2017-val [23] as the non-member set. We randomly select 2500 images from each dataset. We test two scenarios: knowing the ground truth text, denoted as Laion5 vs. COCO; and generating text through BLIP [21], denoted as Laion5 BLIP vs. COCO. For the new dataset, we randomly generated 500 images using DALL-E 3, with prompts derived from the captions associated with each image in LAION-5. This dataset serves as the non-member set, while the corresponding 500 images from LAION-5 serve as the member set.
We use the hyperparameters from the papers [5, 18] to run the baseline methods. For our algorithms ReDiffuse  and ReDiffuse+, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [44] to measure the differences between two images. Other hyperparameters remain the same as in the baseline methods.
@@ Appendix B More Experiment Results
We use different average numbers $n$ and test the influence on the results. Besides the figures of AUC in the main paper, the figures of ASR are also plotted in Figure 8. We observe that in the DDIM setup, averaging the images from multiple independent samples as the output further improves accuracy. In the Stable Diffusion setup, averaging multiple images is not necessary as it does not improve ASR performance.
Figure 8: The impact of average numbers. Left: DDIM model on CIFAR-10. Averaging can further improve the performance of our algorithm. Right: Stable Diffusion model on LAION-5. Averaging does not improve performance.
@@
We adjust the diffusion step $t$ to examine its impact on the results. We plot the graphs for CIFAR100 (Figure 9), STL10 (Figure 10), and Laion5 vs. COCO (Figure 11). From the results, we see that our algorithm can achieve high performance over a wide range of diffusion steps, especially for Stable Diffusion. This opens up possibilities for practical applications in real-world scenarios.
@@
Figure 11: The impact of diffusion steps. We use the Stable Diffusion model with the Laion5 vs. COCO dataset for evaluation and test different diffusion steps for inference. Specifically, using relatively small diffusion steps results in better performance.
Impact of $L_{p}$ norm
For the DDIM model, we train a neural network using the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector v for each image. This corresponds to the $L_{1}$ norm. Then we change the $L_{p}$ norm of our neural network classifier to see if it has any influence on the results. The results are in Figure 12:
Figure 12: The impact of $L_{p}$ norms. Our algorithm remains effective for $p=1,2,3,4$.
From the results, we know that our algorithm is effective for $p=1,2,3,4$. We should not choose very large $p$ to train the neural network.
Impact of Sampling Intervals
We change the sampling intervals to see if there is any influence on the results. In addition to the AUC figures in the main paper, the ASR figures are also plotted in Figure 13. From the results, we observe that the different sampling intervals have minimal impact on ASR.
Figure 13: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Stable Diffusion model on LAION-5. We find that adjusting the sampling interval does not significantly affect the ASR.

### v2

@@ 5 Experiments
In this section, we evaluate the performance of our methods across various datasets and settings. We follow the same experiment setup in previous papers [6, 19]. The detailed experimental settings, including datasets, models, and hyper-parameter settings can be found in Appendix A.
@@ 5.2 MIA with DDIM Models
We follow the experimental setup of [6, 19]. We train a DDIM model on the CIFAR-10/100 [21] and STL10-Unlabeled datasets [2], using the image generation step $T=1000$ and sampling interval $k=100$. For all the datasets, we randomly select 50% of the training samples to train the model and denote them as members. The remaining 50% are utilized as nonmembers. We use [29], [6], [19] as our baseline methods. We fix the diffusion step at $t=200$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$. We will discuss the impact of the diffusion step and the average number in Section 5.5.
Table 1: Comparison of different methods on four datasets for the DDIM model. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method CIFAR10 CIFAR100 STL10
Algorithm U-Net AUC ASR TP AUC ASR TP AUC ASR TP
Loss [29] $\Box$ 0.88 0.82 14.2 0.92 0.84 20.9 0.89 0.82 15.6
SecMI [6] $\Box$ 0.95 0.90 40.7 0.96 0.90 44.9 0.94 0.88 26.9
PIA [19] $\Box$ 0.95 0.89 48.7 0.96 0.90 47.0 0.94 0.87 29.8
PIAN [19] $\Box$ 0.95 0.89 50.4 0.91 0.85 39.2 0.92 0.86 28.5
ReDiffuse $\blacksquare$ 0.96 0.91 40.7 0.98 0.93 48.2 0.96 0.90 31.9
@@
For the difference function $D(x,\hat{x})$, following the setup in [6], we take the pixel-wise absolute value of $x-\hat{x}$ to obtain a difference vector $v$ for each image. Using the ResNet-18 network [10] and denoting it as $f_{R}$, we perform binary classification on these difference vectors. We use 20% of the data as the training set and obtained the label of each difference vector being classified as a member or nonmember. The difference function is obtained by the negated value of the probability outputed by the neural network predicting as member: $D(x,\hat{x})=-f_{R}(v)$.
The result is shown in Table 1. Our method achieves high performance, surpassing several baseline algorithms in most setups, and does not require access to the internal structure of the model. This demonstrates that our algorithm is highly effective and robust.
5.3 MIA with Diffusion Transformers
We train a diffusion transformer model on the ImageNet [3] dataset following the setup of [32]. We randomly select 100,000 images from the ImageNet training set to train the model with resolutions of either $128\times 128$ or $256\times 256$. For the membership inference attack setup, 1000 images are randomly chosen from our training set as the member set, and another 1000 images are randomly selected from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to take the average of the output images as $\hat{x}$.
We use [29], [6], [19] as our baseline methods. Since these work did not study the case of Diffusion Transformers, we integrate their algorithms into the DiT framework for evaluation. For the difference function $D(x,\hat{x})$, following the setup in [6, 19], we take the L2 norm of $x-\hat{x}$ to measure the differences between two image.The results, presented in Table 2, demonstrate that our method outperforms baseline algorithms and does not require access to the Vision Transformer.
Table 2: Comparison of different methods for Diffusion Transformer using the same set of metrics as Table 1. Previous methods require access to the Vision Transformer, whereas our methods do not. We use AUC, ASR, and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method ImageNet 128 $\times$ 128 ImageNet 256 $\times$ 256
Algorithm Transformer AUC ASR TP AUC ASR TP
Loss [29] $\Box$ 0.83 0.76 10.7 0.78 0.70 7.3
SecMI [6] $\Box$ 0.80 0.73 8.3 0.88 0.80 16.3
PIA [19] $\Box$ 0.97 0.92 32.1 0.91 0.85 6.8
PIAN [19] $\Box$ 0.66 0.64 6.2 0.67 0.66 12.8
ReDiffuse $\blacksquare$ 0.98 0.95 44.1 0.97 0.94 47.3
$\Box$: Require the access of Transformer.  $\blacksquare$: Do not require the access of Transformer.
We also plot ROC curves for the DDIM train on CIFAR-10 and the Diffusion Transformer train on ImageNet $256\times 256$ in Figure 3. The curves further demonstrate the effectiveness of our method.
Figure 3: The ROC curves of various setups. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer on ImageNet $256\times 256$. The curves show that our algorithm outperforms the baseline algorithms.
@@ 5.4 MIA with the Stable Diffusion Model
We conduct experiments on the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experiment setup of [6, 19], use the LAION-5B dataset [39] as member and COCO2017-val [24] as non-member. We randomly select 2500 images in each dataset. We test two scenarios: Knowing the ground truth text, which we denote as Laion5; Not knowing the ground truth text and generating text through BLIP [22], which we denote as Laion5 with BLIP.
For the difference function $D(x,\hat{x})$, since the images in these datasets better correlate with human visual perception, we directly use the SSIM metric [47] to measure the differences between two images. The results, presented in Table 3, demonstrate that our methods achieve high accuracy in this setup, outperforming baseline algorithms by approximately 10%. Notably, our methods do not require access to U-Net.
Table 3: Comparison of different methods for Stable Diffusion using the same set of metrics as Table 1. Again, previous methods require access to U-Net, whereas our methods do not. We use AUC, ASR and TP as the metrics, TP refers to the True Positive Rate when the False Positive Rate is 1%.
Method Laion5 Laion5 with BLIP
Algorithm U-Net AUC ASR TP AUC ASR TP
Loss [29] $\Box$ 0.62 0.61 13.2 0.62 0.62 13.3
SecMI [6] $\Box$ 0.70 0.65 19.2 0.71 0.66 19.8
PIA [19] $\Box$ 0.70 0.66 19.7 0.73 0.68 20.2
PIAN [19] $\Box$ 0.56 0.53 4.8 0.55 0.51 4.4
ReDiffuse $\blacksquare$ 0.81 0.75 20.6 0.82 0.75 21.7
@@ 5.5 Ablation Studies
In this section, we alter some experimental parameters to test the robustness of our algorithm. We primarily focus on the ablation study of DDIM and Diffusion Transformer, while the ablation study related to Stable Diffusion is provided in Appendix B.
@@
We test the effect of using different averaging numbers $n$ on the results, as shown in Figure 4 . It can be observed that averaging the images from multiple independent samples to generate the output $\hat{x}$ further improves accuracy. This observation validates the algorithm design intuition discussed in Section 4.2. Additional figures showing the ASR results are presented in Appendix B.
Figure 4: The impact of average numbers. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer model on Imagenet. In both cases, averaging multiple independent samples proves to be effective in further improving the overall performance of our algorithm, which validates the intuition of our algorithm design.
@@
We adjust the diffusion step $t$ to examine its impact on the results. The experiments are conducted using the DDIM model on CIFAR-10 with diffusion steps for inference. The outcomes are presented in Figure 5. Our findings indicate that as long as a moderate step is chosen, the attack performance remains excellent, demonstrating that our algorithm is not sensitive to the choice of $t$. This further underscores the robustness of our algorithm. We also plot the change of diffusion step for other diffusion models and datasets in the Appendix B.
Figure 5: The impact of diffusion steps on DDIM. We train a DDIM model on the CIFAR-10 dataset and use different diffusion steps for inference. We find that high accuracy can be achieved as long as a moderate step number is chosen. This opens up possibilities for practical applications in real-world scenarios.
@@
In DDIM and Diffusion Transformer, the model uses a set of steps denoted by $\tau_{1},\tau_{2},\ldots,\tau_{T}$. It samples each of these steps to create the image. The spacing between these steps is referred to as the sampling interval. We change the sampling interval $k$ and check the influence on the results. As shown in Figure 6, we adjust this parameter for attacks on both DDIM and Diffusion Transformer. We find that our method achieves high AUC values across different sampling intervals, demonstrating that our detection capabilities are not significantly limited by this parameter. Additionally, in Appendix B, we plot the effect of different sampling intervals on ASR and find that the impact is minimal.
@@ 6 Conclusion, Limitations and Future Directions
In this work, we introduce a novel membership inference attack method specifically designed for diffusion models. Our approach only requires access to the variation API of the model, bypassing the need for internal network components such as the U-net. This represents an advancement for commercial diffusion models, which typically restrict internal access. We demonstrate the effectiveness of our approach across various datasets, showing that it achieves high accuracy in identifying whether an image was included in the training dataset. Our algorithm can detect data misuse by the model, representing a step forward in protecting the copyright of artworks.
However, our method has certain limitations, particularly the requirement for a moderate diffusion step $t$ in the variation API. The algorithm’s accuracy declines when the diffusion step is excessively high. As such, we propose our method as an initial step towards black-box MIA, with a more comprehensive solution left for subsequent exploration.
Future work could focus on developing more robust algorithms capable of handling a broader range of diffusion steps. Improving the interpretability of our method and extending it to other generative models are also valuable directions for further research.
@@ Appendix A Datasets, Models and Hyperparameters
We use NVIDIA RTX 6000 graphics cards for all our experiments.
For DDIM, we follow the training hyperparameters of [6] to train a DDIM model on the CIFAR-10/100 [21] and STL10-Unlabeled [2] datasets, using 1000 image generation steps ($T=1000$) and a sampling interval of $k=100$. The training iterations are set to 800,000. For all the datasets, we randomly select 50% of the training samples to train the model and designate them as members. The remaining 50% of the training samples are used as nonmembers.
We use [29, 6, 19] as our baseline methods. We use their official code repositories and apply the optimal hyperparameters from their papers. For our algorithm, we fix the diffusion step at $t=200$ and independently call the variation API 10 times to average the output images as $\hat{x}$.
For the Diffusion Transformers, we train the model on the ImageNet [3] dataset, using 1000 image generation steps ($T=1000$), while for other training hyperparameters, we follow the setup of [32]. We randomly select 100,000 images from the ImageNet training set to train the model at resolutions of either $128\times 128$ or $256\times 256$. For the $128\times 128$ image size, we use 160,000 training iterations. For the $256\times 256$ image size, we use 300,000 training iterations. These numbers of training iterations are chosen to ensure the generation of high-quality images.
For the membership inference attack setup, 1000 images are randomly selected from our training set as the member set, and another 1000 images are randomly chosen from the ImageNet validation set as the non-member set. We fix the diffusion step at $t=150$ and the DDIM step at $k=50$, and independently call the variation API 10 times to average the output images as $\hat{x}$.
We also use [29, 6, 19] as our baseline methods. Since these papers do not study Diffusion Transformers, we adapt their algorithms to the DiT framework for evaluation. We fix the DDIM step at $k=50$ and choose diffusion steps $t\in[50,100,150,200,250,300]$, recording their optimal solutions under these different hyperparameter settings.
For the Stable Diffusion experiments, we use the original Stable Diffusion model, i.e., stable-diffusion-v1-4 provided by Huggingface, without further fine-tuning or modifications. We follow the experimental setup of [6, 19], using an image generation step of $T=1000$ and a sampling interval of $k=10$. We use the LAION-5B dataset [39] as the member set and COCO2017-val [24] as the non-member set. We randomly select 2500 images from each dataset. We test two scenarios: knowing the ground truth text, denoted as Laion; and generating text through BLIP [22], denoted as Laion with BLIP.
We use the hyperparameters from the papers [6, 19] to run the baseline methods. For our algorithms ReDiffuse, we fix the diffusion step at $t=10$ to call the variation API and directly use the SSIM metric [47] to measure the differences between two images. Other hyperparameters remain the same as in the baseline methods.
@@ Appendix B More Experiment Results
We use different average numbers $n$ and test the influence on the results. Besides the figures of AUC in the main paper, the figures of ASR of DDIM and Diffusion Transformer are also plotted in Figure 7. In addition, we plot the figures of AUC and ASR for Stable Diffusion in Figure 8. We observe that in the DDIM and Diffusion Transformer setup, averaging the images from multiple independent samples as the output further improves accuracy. In the stable diffusion setup, since the image size in the dataset is larger (512x512), the reconstructed images are more stable and not influenced by perturbations at specific coordinates. Therefore, averaging multiple images is not necessary.
Figure 7: The impact of average numbers. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer model on Imagenet. Averaging can further improve the performance of our algorithm.
Figure 8: The impact of average numbers on Stable Diffusion. We plot the AUC and ASR metrics, and averaging does not improve performance.
@@
We adjust the diffusion step $t$ to examine its impact on the results. We train the DDIM model on CIFAR100 (Figure 9), STL10 (Figure 10) dataset, Diffusion Transformer on the Imagenet $256\times 256$ (Figure 11)dataset and Stable Diffusion on Laion5 dataset (Figure 12). From the results, we see that our algorithm can achieve high performance over a wide range of diffusion steps. This opens up possibilities for practical applications in real-world scenarios.
@@
Figure 11: The impact of diffusion steps. We train a Diffusion Transformer model on the Imagenet $256\times 256$ dataset and use different diffusion step for inference. The robust results imply that our algorithm is also not very sensitive to the choice of $t$ in this setup.
Figure 12: The impact of diffusion steps. We use the Stable Diffusion model with the Laion5 dataset for evaluation and test different diffusion steps for inference. Specifically, using relatively small diffusion steps results in better performance.
The Impact of Sampling Intervals
We change the sampling intervals to see if there is any influence on the results. In addition to the AUC figures in the main paper, the ASR figures of DDIM and Diffusion Transformer are also plotted in Figure 13. We also plot the AUC and ASR for Stable Diffusion in Figure 14. From the results, we observe that our algorithm consistently performs well across different sampling intervals.
Figure 13: The impact of sampling intervals. Left: DDIM model on CIFAR-10. Right: Diffusion Transformer on Imagenet. We find that adjusting the sampling interval does not significantly affect of the accuracy.
Figure 14: The impact of sampling intervals on Stable Diffusion. We plot the AUC and ASR metrics, observe that different sampling intervals have minimal impact on it.
The ROC curves
Besides the ROC curves in the main paper, we also plot ROC curves for the Diffusion Transformer train on ImageNet $128\times 128$ and the Stable Diffusion train on Laion5 in Figure 15. The curves further demonstrate the effectiveness of our method.
Figure 15: The ROC curves of various setups. Left: Diffusion Transformer on ImageNet $128\times 128$. Right: The Stable Diffusion on Laion5. The curves show that our algorithm outperforms the baseline algorithms.